# NeuralOM: Neural Ocean Model for Subseasonal-to-Seasonal Simulation

### v4

@@ B.1    Multi-scale Graph Encoder
where, $\mathrm{ESMLP}(\cdot)$ is the Edge Sum MLP (pfaff2020learning). The mesh node features are updated by an MLP:
@@ C.1    Evaluation Metrics
where $\mathbf{A}_{i,j,t}^{\mathcal{K}}$ represents the value of variable $\mathcal{J}$ at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i’=1}^{N_{\text{lat}}}\cos\phi_{i’}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from multiple years of training data. RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for each variable $\mathcal{J}$ at a given lead time $\Delta t$.
@@ C.2    Model Training
For the progressive residual correction framework, we apply Q sub-models to progressively capture subtle changes in ocean variables. We train the first sub-model with M steps finetune. For other sub-models, we train them with N steps finetune. In our experiment, we set Q=2, M=6, and N=10 as an example. Specifically, we train the first sub-model, denoted as $\mathcal{F}^{1}_{1}$ with 200 epochs using 1-step supervision. The learning rate is set to 1e-3. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. After that, we finetune $\mathcal{F}^{1}_{1}$ with 10 epochs using 2-step supervision to get $\mathcal{F}^{2}_{1}$. Then, we finetune $\mathcal{F}^{2}_{1}$ with 10 epochs using 3-step supervision to get $\mathcal{F}^{3}_{1}$. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. The model $\mathcal{F}^{4}_{1}$, $\mathcal{F}^{5}_{1}$, and $\mathcal{F}^{M}_{1}$ are finetuned same as $\mathcal{F}^{3}_{1}$. Then we freeze the parameters of $\mathcal{F}^{M}_{1}$ and train the second sub-model $\mathcal{F}^{1}_{2}$ with 200 epochs. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. In practice, the model converges after a few epochs, and we stop the training early. The the $\mathcal{F}^{N}_{2}$ can be get step by step with same training produce as $\mathcal{F}^{1}_{2}$. Furthermore, the Q-th model $\mathcal{F}^{N}_{Q}$ can be trained step by step if a higher accuarcy is required. We train all baseline models and our NeuralOM using the same training framework. We will open-source all of our materials—including all training codes, detailed training procedures, testing codes, pre-trained weights, training logs, and more.

### v5

@@ B.1    Multi-scale Graph Encoder
where, $\mathrm{ESMLP}(\cdot)$ is the Edge Sum MLP (Pfaff et al. 2020). The mesh node features are updated by an MLP:
@@ C.1    Evaluation Metrics
where $\mathbf{A}_{i,j,t}^{\mathcal{K}}$ represents the value of variable $\mathcal{J}$ at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i’=1}^{N_{\text{lat}}}\cos\phi_{i’}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from multiple years of training data. RMSE and ACC are averaged across all time steps and over non-NaN spatial grid points, providing summary statistics for each variable $\mathcal{J}$ at a given lead time $\Delta t$.
@@ C.2    Model Training
For the progressive residual correction framework, we apply Q sub-models to progressively capture subtle changes in ocean variables. We train the first sub-model with M steps finetune. For other sub-models, we train them with N steps finetune. In our experiment, we set Q=2, M=6, and N=10 as an example. Specifically, we train the first sub-model, denoted as $\mathcal{F}^{1}_{1}$ with 200 epochs using 1-step supervision. The learning rate is set to 1e-3. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. After that, we finetune $\mathcal{F}^{1}_{1}$ with 10 epochs using 2-step supervision to get $\mathcal{F}^{2}_{1}$. Then, we finetune $\mathcal{F}^{2}_{1}$ with 10 epochs using 3-step supervision to get $\mathcal{F}^{3}_{1}$. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. The model $\mathcal{F}^{4}_{1}$, $\mathcal{F}^{5}_{1}$, and $\mathcal{F}^{M}_{1}$ are finetuned same as $\mathcal{F}^{3}_{1}$. Then we freeze the parameters of $\mathcal{F}^{M}_{1}$, and the second sub-model $\mathcal{F}^{N}_{2}$ can be trained in two ways. One way is that we first pre-train $\mathcal{F}^{1}_{2}$, and then the $\mathcal{F}^{N}_{2}$ can be get step by step. The other way is that we directly train $\mathcal{F}^{N}_{2}$. In practice, the model converges after a few epochs, and we stop the training early. Furthermore, the Q-th model $\mathcal{F}^{N}_{Q}$ can be trained step by step if a higher accuarcy is required. We train all baseline models and our NeuralOM using the same training framework. We will open-source all of our materials—including all training codes, detailed training procedures, testing codes, pre-trained weights, etc.